1. A project has and initial cost of $52,000, expected net cash inflows of $11,000 per year for 7 years, and a cost of capital of 9%.

a. What is the project’s NPV?

b. What is the project’s IRR?

c. What is the projects payback period?
Answer #1

a.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=11000[1-(1.09)^-7]/0.09

=11000*5.032952835

=55362.48

NPV=Present value of inflows-Present value of outflows

=55362.48-$52000

=3362.48(Approx).

2.

Let irr be x%
At irr,present value of inflows=present value of outflows.

52,000=11,000/1.0x+11,000/1.0x^2+............+11,000/1.0x^7

Hence x=irr=10.90%(Approx).

3.Payback period=initial cost/annual cash flows

=(52,000/11,000)

=4.73 years(Approx).